the sum of the interior angles is 5400 degrees.
how many sides does the polygon have?

Found 2 solutions by Alan3354, jeseca_1964:
Answer by Alan3354(69443)

180*(n-2) = 5400

Answer by jeseca_1964(5)

180(n-2) = 5400
180(n-2/180 = 5400/180
n-2 = 30
n = 30 + 2
n = 32